OpenStudy (anonymous):

HOW DO I FIND THE DISCRIMINANT OF 2X^2+11X-21=0 ?

Directrix (directrix):

See formula attached.

Directrix (directrix):

11^2 -4*2*(-21)= ? @stupidfresh305

OpenStudy (jhannybean):

and when the discriminant, \(b^2 -4ac \gt 0\) , you have 2 solutions, \(b^2-4ac < 0\), you have 0 solutions, while \(b^2 -4ac = 0\) would render 1 solution :)

OpenStudy (calculusfunctions):

Do you still need help with this question? Because you just have to do what @Jhannybean suggested.

Directrix (directrix):

The question is to give the discriminant. The nature of the roots as determined by the discriminat is not part of the question as I read it.

OpenStudy (anonymous):

Simplifying 2x2 + -11x + -21 = 0 Reorder the terms: -21 + -11x + 2x2 = 0 Solving -21 + -11x + 2x2 = 0 Solving for variable 'x'. Factor a trinomial. (-3 + -2x)(7 + -1x) = 0 Subproblem 1 Set the factor '(-3 + -2x)' equal to zero and attempt to solve: Simplifying -3 + -2x = 0 Solving -3 + -2x = 0 Move all terms containing x to the left, all other terms to the right. Add '3' to each side of the equation. -3 + 3 + -2x = 0 + 3 Combine like terms: -3 + 3 = 0 0 + -2x = 0 + 3 -2x = 0 + 3 Combine like terms: 0 + 3 = 3 -2x = 3 Divide each side by '-2'. x = -1.5 Simplifying x = -1.5 Subproblem 2 Set the factor '(7 + -1x)' equal to zero and attempt to solve: Simplifying 7 + -1x = 0 Solving 7 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-7' to each side of the equation. 7 + -7 + -1x = 0 + -7 Combine like terms: 7 + -7 = 0 0 + -1x = 0 + -7 -1x = 0 + -7 Combine like terms: 0 + -7 = -7 -1x = -7 Divide each side by '-1'. x = 7 Simplifying x = 7